OpenStudy (agentjamesbond007):

[Related Rates] If V is the volume of a cube with edge length x and the cube expands as time passes, find dv/dt in terms of dx/dt.

OpenStudy (anonymous):

\[\frac{dv}{dt}=\frac{dv}{dx}\times \frac{dx}{dt}\]

OpenStudy (anonymous):

and since \[v=x^3\] you get \[\frac{dv}{dt}=3x^2\frac{dx}{dt}\]